1) You are given the following information from a well known corporation. Conduct an analysis of the data. This analysis should include year-over-year changes in terms of percentage and absolute/net for:

  • Revenues
  • Operating Income
  • Net Income

Calculate the following ratios for each year:

Net Profit Margin

Operating Profit Margin

Net Return on Total Assets (ROA)

Return on Stockholders Equity

Here's the data:

                                                 2017          2016            2015         2014          2013

Sales (Sales Revenue)                12917         11699     10108         9006           8032

Operating Income                               883              743              548          438             284

Profit After Taxes                        551              466              343          246             147

Total Assets                            5538            5294        4292        3987      3783  

Total Stockholders Equity     3878          3802         2991        2373          1628

Be sure include a summary on your assessment. What was the trend with Revenues? Operating Income? Net Income? Using the Appendix, did the ratios meet the criteria for good performance? Is the trend positive, negative, about the same?

You must show the work! Present your completed problem demonstrating the analysis in a table or chart format!
Answer #1
Details/ Year 2017 2016 2015 2014 2013
Sales (Sales Revenue)     12,917        11,699        10,108        9,006        8,032
Operating Income          883              743              548           438           284
Profit After Taxes          551              466              343           246           147
Total Assets       5,538          5,294          4,292        3,987        3,783
Total Stockholders Equity       3,878          3,802          2,991        2,373        1,628
a) Ratios:
Net profit margin = Profit after tax/Sales , for example 2017 = 551/12917% = 4.27%, similarly for other years 4.27% 3.98% 3.39% 2.73% 1.83%
Operating profit margin = Operating income/Sales , for example 2017 = 883/12917% = 6.84%
similarly for other years
6.84% 6.35% 5.42% 4.86% 3.54%
Net Return on Total Assets (ROA) = Profit after tax/Total assets , for example 2017 = 551/5538% = 6.84%, similarly for other years 9.95% 8.80% 7.99% 6.17% 3.89%
Return on stockholders equity = Profit after tax/Total stockholders equity, for example 2017 = 551/3878% = 6.84%, similarly for other years 14.21% 12.26% 11.47% 10.37% 9.03%
b) Year on Year changes 2017 2016 2015 2014
Revenues:
Revenues (Absolute) - calculated as change between two years, for example 2017 is revenue of 2017 - revenue of 2016 which is 12917-11699=1218, similarly for other years       1,218          1,591          1,102           974
Revenues (in %) - calculated as (change between two years)/previous year's revenue, for example 2017 is = 1218/11699% = 10.41% 10.41% 15.74% 12.24% 12.13%
Assessment:
Revenues has been increasing on each year on absolute terms as well as on percentage terms, however there is a slight decreasing in 2017 in terms of both absolute and percentage terms as compared to previous year for which explanation should be sought for.
Similarly calculate change in operating income and net income as done above
Operating income:
Operating income (Absolute) - calculated as change between two years, for example 2017 is 883 - 743 which is 140, similarly for other years          140              195              110           154
Operating income (in %) - calculated as (change between two years)/previous year's operating income, for example 2017 change is calculated above = 140/743% = 18.84% 18.84% 35.58% 25.11% 54.23%
Assessment:
Operating income has a positive absolute number and positive percentage as well, however the there is an erratic trend, it is not flowing a regular trend and has been fluctuating though remaining positive throughout the years.
Net income:
Net income/profit after tax (Absolute) - calculated as change between two years, for example 2017 is 551-466 which is 85, similarly for other years            85              123                97              99
Net income (in %) - calculated as (change between two years)/previous year's operating income, for example 2017 change is calculated above = 85/466% = 18.24% 18.24% 35.86% 39.43% 67.35%
Assessment:
Net income has a positive absolute number and positive percentage as well, however the net income in percentage has been showing a declining trend as compared to previous years for which an explanation can be sought for.